How many mols of gas are present in a sample of oxygen if it is found to be held at 4.98 atm in a 29.5 L container at 29.0 C ?
1
Expert's answer
2020-04-06T11:57:53-0400

How many mols of gas are present in a sample of oxygen if it is found to be held at 4.98 atm in a 29.5 L container at 29.0 C ?


Solution:

PV=nRT

"n= \\frac{PV}{RT}=\\frac{4.98*29.5}{0.0821(273+29)}=5.925 moles\n\n\u200b\t\n\n\u200b"

Answer = 5.925 moles or approximate 6 moles